Dumoor Appliance Center sells and services several brands of major appliances. Past sales for a particular...

  1. Dumoor Appliance Center sells and services several brands of major appliances. Past sales for a particular model of refrigerator have resulted in the following probability distribution for demand:

Demand per week

0

1

2

3

Probability

0.10

0.40

0.40

0.10

The lead time, in weeks, is described by the following distribution:

Lead Time (weeks)

1

2

3

Probability

0.15

0.80

0.05

Based on the cost considerations as well as storage space, the company has decided to order 6 of these each time an order is placed. The holding cost is $2 per week for each unit that is left in inventory at the end of the week. The stockout cost has been set at $30 per stockout. The company has decided to place an order whenever there are only 2 refrigerators left at the end of the week. Simulate 10 weeks of operation for Dumoor Appliance assuming there are currently 5 units in inventory. Determine what the weekly stockout cost and weekly holding cost would be for the problem. (10 points)
Answer #1

The simulation model is following:

F G H K L M Lead Time 1 Prob. 0.15 0.80 0.05 Random Number Range 0 15 16 95 96 100 2 3 Order Quantity Reorder Point 6 2 A B с

EXCEL FORMULA:

А B с D E G H K Prob. Random Number Range 0.15 0.8 0.05 0 =J2+1 =J3+1 =H2*100+12 =H3*100+13-1 =H4*100+14-1 4 6 2 2 1 Demand P

D H 1 Prob. Random Number Range 0.15 0.8 0.05 0 =J2+1 =J3+1 =H2*100+12 =H3*100+13-1 =H4*100+14-1 6 2 30 Random Number Range L